Jim rides his skateboard (starting at rest) down a ramp that is in the shape of a quarter circle with radius of 3.9 m. At the bottom of the ramp, Jim is moving at 7.5 m/s. Jim and his skateboard have a mass of 66.9 kg. How much work is done by friction as the skateboard goes down the ramp?
Answer #1

work done by friction = PE - KE(At bottom)

wf = m*g*r - 1/2 * m*v^2

wf = 66.9 * 9.8 * 3.9 - 1/2 * 66.9 * (7.5)^2

wf = 675.3555